Wednesday 11 August 1999 7.00am EST
|The following is an overview of the events that occurred over the last 24 hours
Completed Gyro Low Rate Bias Determination
Continued ACIS Background Measurements
Optical Cavity pressure measurement reading was 7.8E-06 compared with 8.1, which had been holding at 8.1 for the past 48 hours.
Complete Earth Scan #24.
Events scheduled for the next 24 hours
Return to Status Reports